# Gaussian Pancakes: Geometrically-Regularized 3D Gaussian Splatting for Realistic Endoscopic Reconstruction

### v1

@@ Abstract
[…]
@@ 1 Introduction
Colorectal cancer remains a major global health challenge, consistently ranking among the top three cancers in prevalence and mortality [WorldHealthOrganization_2022]. The slow progression of colorectal cancer offers detecting pre-cancerous growth through colonoscopy. Despite several AI-based methods proposed in the literature for polyp detection during colonoscopy [barua2020artificial], challenges persist, particularly in identifying polyps hidden behind folds, which account for up to three-quarters of all missed cases [pickhardt2004location]. Colonoscopists encounter difficulties in thorough surface inspection, due to the use of a monocular camera with limited field of view and lack of 3D information while navigating the colon’s complex structure. Generating full 3D reconstructions with high quality textures from endoscopic images in near real-time would enable improved diagnosis and treatment through downstream tasks like automated surgical path planning [9712211], AR/VR training environments[sun2023dynamic], AI-based polyp detection [barua2020artificial], and for determining missing regions for re-inspection [ma2021rnnslam].
@@
In response to these limitations, we propose Gaussian Pancakes, an approach that synergizes a geometrically-regularized 3D Gaussian Splatting with a colonscopy-tailored Simultaneous Localization and Mapping (SLAM) system, RNNSLAM [ma2021rnnslam], to enhance texture rendering and anatomical accuracy while boosting training and rendering speeds. Our method capitalizes on the strengths of RNNSLAM, which generates camera poses, depth maps, and a rudimentary surface reconstruction from only the monocular video inputs in real time, providing a robust alternative to traditional Structure-from-Motion (SfM) methods. Although RNNSLAM alone yields surface reconstructions, the lack of photorealism and anatomical detail limits its use in novel view synthesis tasks like AI-based polyp detection. Our proposed Gaussian Pancakes technique overcomes these drawbacks by replacing surfels with 3D Gaussians and incorporating surface-based regularization, resulting in significantly accurate radiance fields and vivid textures. Our key contributions are as follows: 1) SLAM Integration: We created a pipeline that merges 3D GS with SLAM for photorealistic 3D colonoscopic reconstructions, that demonstrated robustness over using a SfM method seen in Fig. 1 (C) and Fig. 1 (D). 2) Improvement of Base GS Method: We improved the 3D GS method by incorporating geometric and depth regularizations, which aligns Gaussians with the surfaces by effectively “pancaking” them. This enhancement reduces geometric error and artifacts in novel view synthesis, seen in Fig. 1. 3) Endoscopic Radiance Field Advancements: We optimized the training and rendering process for radiance fields in surgical scenes, reducing training times to roughly 2 minutes and improving image rendering speeds by more than 100 times, making it vastly more applicable in clinical settings.
@@ 2 Method
Gaussian Pancakes employs 3D GS for realistic texture rendering and RNNSLAM for 3D reconstruction. As shown in Fig . 2, it starts with RNNSLAM generated camera poses, depth maps, and meshes. This mesh is sampled, using stratified sampling, to create a sparse RGB point cloud [alexa2004stratified]. Around these points, Gaussians with corresponding attributes are initialized (Sec. 2.2). Supervision is twofold: photometric and geometric regularization, with the addition of geometric regularization ensuring Gaussians align with the surface and limit floating artifacts.
@@ 2.1 Preliminaries of RNNSLAM
RNNSLAM [ma2021rnnslam] uses sparse depth estimation from a SfM method as a proxy for ground truth to train a recurrent neural network for depth and camera pose prediction. Bundle-adjusted direct sparse odometry is used to jointly optimize the predicted poses and depth by minimizing the intensity difference over a window of recent frames. The fusion pipeline is then used to reconstruct colon meshes with textures, enabling the detection of missing regions. While this approach has shown effectiveness in reconstructing dense 3D colon maps from visible sections in real colonoscopy, RNNSLAM suffers from pose estimation drift, resulting in texture misalignment and structural collapse in the fused meshes, which is a common problem among similar methods [rau2023simcol3d]. Our proposed method uses the refined depth maps, camera poses and sparsely sampled points on converged meshes of RNNSLAM to provide high quality texture renderings of the structures.
Figure 2: Proposed Gaussian Pancakes’ pipeline highlighting our contributions: A) RNNSLAM for mesh, camera poses, and depth maps; B) 3D GS initialization, and C) Geometric and depth regularizations, distinguishing our approach from traditional 3D GS.
@@ 2.2 Preliminaries 3D Gaussian Splatting
3D GS employs Gaussian distributions for scene representation and leverages GPU-accelerated rendering for efficient optimization [kerbl20233d], [lassner2021pulsar]. 3D GS method models scenes using 3D Gaussians around the input RGB point cloud, represented by position $\bm{\mu}$, scale $s$, orientation as quaternion $\mathbf{q}$, spherical coefficients $SH$, and opacity $\sigma$. Each Gaussian has a covariance matrix $\Sigma$ describing the skew and variance of the distribution in 3 dimensions, calculated using the scaling matrix $S$ and rotation matrix $R$ by
@@
A Gaussian $G(\mathbf{x})$ can be described by the multivariate Gaussian probability density:
@@
These Gaussians are expressed in 3D space but are projected onto 2D for image rendering given by a viewing transformation $W$ and the covariance matrix $\Sigma$. The covariance matrix in the camera coordinate system is given by $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$, where $J$ is the Jacobian matrix of the affine approximation of $W$. The splatting process, integral to 3D GS, involves differentiable point-based alpha blending. The final image pixel color is accumulated from contributions along a ray, considering the densities and transmittances. The color of each Gaussian is represented using spherical harmonic coefficients, capturing view-dependent effects [ramamoorthi2001efficient]. The Gaussian attributes are optimized using the Adam optimizer [kingma2017adam] through a rendering process, combined with steps to adjust Gaussian density.
@@ 2.3 Proposed Gaussian Pancakes: Extending 3D GS for Endoscopy
Instead of relying on the sparse point cloud output from an SfM method, our approach utilizes RNNSLAM for pose estimation, depth maps, and RGB point cloud generation [schoenberger2016sfm]. Directly using the SfM output yielded less than ideal results when paired with 3D GS, failing to capture any realistic geometry as evidenced from Fig. 1. This substitution addresses SfM’s limitations in low-variation environments where uniform textures can impede accurate feature matching and depth perception, thoroughly described in the RNNSLAM paper [ma2021rnnslam].
@@
[…]
@@
Geometric regularization: Despite the introduction of the depth loss, challenges persist due to the lack of surface constraints. Artifacts which present as floating objects often emerge as the rendered view gets further away from original camera positions. To mitigate this, we incorporate a geometric loss that constrains the Gaussians in accordance with the surface’s principal curvature, effectively “pancaking” the Gaussians along the surface. For a point cloud $P=\{\mathbf{p}_{m}\}$ where $\{\mathbf{p}_{m}\}=(x_{m},y_{m},z_{m})$ represents the $m$-th point, we compute the normal vector $\mathbf{n}_{p}$ for each query point $\mathbf{p}_{m}$ in $P$. This computation proceeds under the premise that the local neighborhood of any point can be approximated by a tangent plane, inferred from its 10 nearest neighbors. We denote these neighbors as $\{\mathbf{n}_{j}\}_{j=1}^{k}$, where $k$ = 10. Then centering the nearest neighbors by $\mathbf{n^{\prime}}_{j}=\mathbf{n}_{j}-\frac{1}{k}\sum_{j=1}^{k}\mathbf{n}_{j}$. The covariance of $\mathbf{n^{\prime}}$, $C$, is calculated by
@@
Eigen decomposition of $C$ yields eigenvalues and corresponding eigenvectors $\mathbf{v}_{i}$. The normal vector, $\mathbf{n}_{p}$, for each query point, $\mathbf{p}_{m}$, is taken as the eigenvector corresponding to the smallest eigenvalue: $\mathbf{n_{p}}=\mathbf{v_{1}}$. The concatenated matrix of normal vectors and their corresponding points, $N$, is indexed using a GPU-accelerated index structure, enhancing search efficiency [johnson2017billionscale].
Normal vectors for each 3D Gaussian are efficiently recalculated for new Gaussians per iteration using a custom CUDA kernel to accommodate the dynamic scene efficiently. For each Gaussian, we construct its rotation matrix $R$ and identify the axis of least spread by finding the minimum index of its scale vector $s$. This axis, selected from the standard Cartesian set $e_{x},e_{y},e_{z}$} as $\mathbf{e}_{min}$, serves as the primary basis vector for computing the normal vector $\mathbf{n}_{GP}=R\mathbf{e}_{min}$. Utilizing normals $A$ from the closest point cloud points and Gaussian normals $B$, we calculate the geometric loss using the cosine similarity as follows:
[…]
@@
with $\mathcal{L}_{Image}$ an L1 loss and $\mathcal{L}_{D-SSIM}$, a structural similarity term, between rendered and ground truth images. This loss optimizes the trade-offs between visual fidelity, depth error, and geometric alignment with empirically set weight parameters.
@@ 3 Experiments
3.1 Dataset and Evaluation Metrics
Our evaluation encompasses three distinct datasets: Simulation [zhang2020template], In-Vivo [ma2021rnnslam], and Phantom [bobrow2023colonoscopy]. The Simulation dataset, featuring sequences of the cecum, rectum, and sigmoid regions at a resolution of 320 $\times$ 240, is derived from software simulations in Unity using the graphics engine[zhang2020template] for image rendering. Depth and pose references, however, are obtained from RNNSLAM. The In-Vivo dataset comprises real colonoscopy sequences each at 270 $\times$ 216 resolution from the RNNSLAM study[ma2021rnnslam]. We selected three distinct sequences ("cecum_t4_b", "desc_t4_a", and "transt_t1_a") covering the cecum, descending colon and transcending sections from the publicly-available Phantom dataset (C3VD[bobrow2023colonoscopy]), featuring wide-angle colonoscopy videos at 1350 $\times$ 1080 resolution. C3VD’s simple camera trajectories caused large gaps in reconstructions due to insufficient scene coverage for RNNSLAM’s mesh surface fusion step, limiting us to choose three datasets with satisfactory RNNSLAM performance to effectively evaluate 3D renderings.
@@ 3.1 Dataset and Evaluation Metrics
3.2 Implementation
For a fair comparison with REIM-NeRF and NeRF, which require extensive training time and increase training time with higher resolutions, we adjusted the resolution of the phantom dataset (C3VD) to 338 $\times$ 270, decreased similar to that reported by REIM-NeRF, while maintaining original resolutions for other datasets due to their manageable sizes. Our methodology initiates with RNNSLAM, rendering 10 FPS on an Intel i7 single core CPU and Nvidia GeForce GTX 1080 GPU. Gaussian Pancakes uses inferred depth maps for supervision, along with inferred camera poses and points for initial setup, across all training sequences. Subsequently, Gaussian Pancakes models trained following the Kerbl et al. [kerbl20233d] paper’s hyperparameters, but are trained for 7,000 iterations instead of 30,000, densifying Gaussians until the 4,000th iteration. We set $\delta$ at 0.2 in Eq. 3, with hyperparameters in Eq. 6 $\lambda_{1},\lambda_{2},\text{and }\lambda_{3}$ finely tuned to 0.2, 0.6, and 0.2 (selected empirically), respectively, for a balance of photometric, reconstruction, and geometric results. To facilitate non-uniform scaling, geometric loss integration commences after the first 1000 iterations. All models are trained for 1-2 minutes on an NVIDIA RTX A6000.
Model Dataset PSNR ↑

### v2

@@ Abstract
Within colorectal cancer diagnostics, conventional colonoscopy techniques face critical limitations, including a limited field of view and a lack of depth information, which can impede the detection of precancerous lesions. Current methods struggle to provide comprehensive and accurate 3D reconstructions of the colonic surface which can help minimize the missing regions and reinspection for pre-cancerous polyps. Addressing this, we introduce “Gaussian Pancakes”, a method that leverages 3D Gaussian Splatting (3D GS) combined with a Recurrent Neural Network-based Simultaneous Localization and Mapping (RNNSLAM) system. By introducing geometric and depth regularization into the 3D GS framework, our approach ensures more accurate alignment of Gaussians with the colon surface, resulting in smoother 3D reconstructions with novel viewing of detailed textures and structures. Evaluations across three diverse datasets show that Gaussian Pancakes enhances novel view synthesis quality, surpassing current leading methods with a 18% boost in PSNR and a 16% improvement in SSIM. It also delivers over 100$\times$ faster rendering and more than 10$\times$ shorter training times, making it a practical tool for real-time applications. Hence, this holds promise for achieving clinical translation for better detection and diagnosis of colorectal cancer. Code: https://github.com/smbonilla/GaussianPancakes.
Keywords:
Gaussian Splatting 3D Reconstruction Robotic Surgery
@@ 1 Introduction
Colorectal cancer remains a major global health challenge, consistently ranking among the top three cancers in prevalence and mortality [WorldHealthOrganization_2022]. The slow progression of colorectal cancer offers detecting pre-cancerous growth through colonoscopy. Despite several AI-based methods proposed in the literature for polyp detection during colonoscopy [barua2020artificial], challenges persist, particularly in identifying polyps hidden behind folds, which account for up to three-quarters of all missed cases [pickhardt2004location]. Colonoscopists encounter difficulties in thorough surface inspection, due to the use of a monocular camera with limited field of view and lack of 3D information while navigating the colon’s complex structure. Generating full 3D reconstructions with high-quality textures from endoscopic images in near real-time would enable improved diagnosis and treatment through downstream tasks like automated surgical path planning [9712211], AR/VR training environments[sun2023dynamic], AI-based polyp detection [barua2020artificial], and for determining missing regions for re-inspection [ma2021rnnslam].
@@
In response to these limitations, we propose Gaussian Pancakes, an approach that synergizes a geometrically-regularized 3D GS with a colonoscopy-tailored Simultaneous Localization and Mapping (SLAM) system, RNNSLAM [ma2021rnnslam], to enhance texture rendering and anatomical accuracy while boosting training and rendering speeds. Our method capitalizes on the strengths of RNNSLAM, which generates camera poses, depth maps, and rudimentary surface reconstruction from only the monocular video inputs in real-time, providing a robust alternative to traditional Structure-from-Motion (SfM) methods. Although RNNSLAM alone yields surface reconstructions, the lack of photorealism and anatomical detail limits its use in novel view synthesis tasks. Gaussian Pancakes overcomes these drawbacks by replacing surfels with 3D Gaussians and incorporating surface-based regularization, resulting in significantly accurate radiance fields and vivid textures. Our key contributions are: 1) A pipeline that integrates 3D GS with SLAM for photorealistic 3D colonoscopic reconstructions, demonstrating robustness over using a SfM method seen in Fig. 1 (C-D); 2) Improved 3D GS method by incorporating geometric and depth regularizations, aligning Gaussians with the surfaces by effectively “pancaking” them. This enhancement reduces geometric error and artifacts in novel view synthesis, seen in Fig. 1; 3) Optimized training and rendering process for radiance fields in surgical scenes, reducing training times to roughly 2 minutes and improving image rendering speeds by more than 100 times, making it more applicable in clinical settings.
@@ 2 Method
Gaussian Pancakes employs 3D GS for realistic texture rendering and RNNSLAM for 3D reconstruction. As shown in Fig . 2, it starts with RNNSLAM-generated camera poses, depth maps, and meshes. This mesh is sampled, using stratified sampling, to create a sparse RGB point cloud [alexa2004stratified]. Around these points, Gaussians with corresponding attributes are initialized (Sec. 2.2). Supervision is twofold: photometric and geometric regularization, with the addition of geometric regularization ensuring Gaussians align with the surface and limit floating artifacts.
@@ 2.1 Preliminaries of RNNSLAM
RNNSLAM [ma2021rnnslam] uses sparse depth estimation from a SfM method as a proxy for ground truth to train a recurrent neural network for depth and camera pose prediction. Bundle-adjusted direct sparse odometry is used to jointly optimize the predicted poses and depth by minimizing the intensity difference over a window of recent frames. The fusion pipeline is then used to reconstruct colon meshes with textures, enabling the detection of missing regions. While this approach has shown effectiveness in reconstructing dense 3D colon maps from visible sections in real colonoscopy, RNNSLAM suffers from pose estimation drift, resulting in texture misalignment and structural collapse in the fused meshes, which is a common problem among similar methods [rau2023simcol3d]. Our proposed method uses the refined depth maps, camera poses and sparsely sampled points on converged meshes of RNNSLAM to provide high-quality texture renderings of the structures.
Figure 2: Proposed Gaussian Pancakes’ pipeline highlighting our contributions: A) RNNSLAM for mesh, camera poses, & depth maps; B) 3D GS initialization, C) Geometric & depth regularizations, distinguishing our approach from traditional 3D GS.
@@ 2.2 Preliminaries 3D Gaussian Splatting
3D GS employs Gaussian distributions for scene representation and leverages GPU-accelerated rendering for efficient optimization [kerbl20233d], [lassner2021pulsar], [chen2024survey]. 3D GS method models scenes using 3D Gaussians around the input RGB point cloud, represented by position $\bm{\mu}$, scale $s$, orientation as quaternion $\mathbf{q}$, spherical coefficients $SH$, and opacity $\sigma$. Each Gaussian has a covariance matrix $\Sigma$ describing the skew and variance of the distribution in 3 dimensions, calculated using the scaling matrix $S$ and rotation matrix $R$ by
@@
A Gaussian can be described by the multivariate Gaussian probability density:
@@
These Gaussians are expressed in 3D space but are projected onto 2D for image rendering given by a viewing transformation $W$ and the covariance matrix $\Sigma$. The covariance matrix in the camera coordinate system is given by $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$, where $J$ is the Jacobian matrix of the affine approximation of $W$. The splatting process involves differentiable point-based alpha blending. The final image pixel color is accumulated from contributions along a ray, considering the densities and transmittances. The color of each Gaussian is represented using spherical harmonic coefficients, capturing view-dependent effects [ramamoorthi2001efficient]. The Gaussian attributes are optimized using the Adam optimizer [kingma2017adam] through a rendering process, combined with steps to adjust Gaussian density.
@@ 2.3 Proposed Gaussian Pancakes: Extending 3D GS for Endoscopy
Instead of relying on the sparse point cloud output from an SfM method, our approach utilizes RNNSLAM for pose estimation, depth maps, and RGB point cloud generation [schoenberger2016sfm]. This substitution addresses SfM’s limitations in low-variation environments where uniform textures can impede accurate feature matching and depth perception, thoroughly described in the RNNSLAM paper [ma2021rnnslam].
@@
$\mathcal{L}_{D}(i)=\left\{\begin{array}[]{lr}0.5\Delta D_{i}^{2},&\text{ if }|\Delta D_{i}|<\delta\\ \delta(\Delta D_{i}-0.5\delta),&\text{otherwise }\end{array}\right\}$ (3)
@@
Geometric regularization: Despite the depth loss, challenges persist due to the lack of surface constraints. Artifacts which present as floating objects often emerge as the rendered view gets further away from original camera positions. To mitigate this, we incorporate a geometric loss that constrains the Gaussians in accordance with the surface’s principal curvature, effectively “pancaking” the Gaussians along the surface. For a point cloud $P=\{\mathbf{p}_{m}\}$ where $\{\mathbf{p}_{m}\}=(x_{m},y_{m},z_{m})$ represents the $m$-th point, we compute the normal vector $\mathbf{n}_{p}$ for each point $\mathbf{p}_{m}$ in $P$. This computation assumes the local neighborhood of any point can be approximated by a tangent plane inferred from its 10 nearest neighbors. We denote these neighbors as $\{\mathbf{n}_{j}\}_{j=1}^{k}$, where $k$ = 10. Centering the nearest neighbors by $\mathbf{n^{\prime}}_{j}=\mathbf{n}_{j}-\frac{1}{k}\sum_{j=1}^{k}\mathbf{n}_{j}$. The covariance of $\mathbf{n^{\prime}}$, $C$, is calculated
@@
Eigen decomposition of $C$ yields eigenvalues and eigenvectors $\mathbf{v}_{i}$. The normal vector, $\mathbf{n}_{p}$, for each point, $\mathbf{p}_{m}$, is the eigenvector corresponding to the smallest eigenvalue. The concatenated matrix of normal vectors and their points, $N$, is indexed using a GPU-accelerated structure, enhancing search efficiency [johnson2017billionscale].
Normal vectors for each 3D Gaussian are efficiently recalculated per iteration with a custom CUDA kernel for the dynamic scene. For each Gaussian’s rotation matrix $R$, we find the axis of least spread by locating the minimum index of its scale vector $s$. This axis, selected from the standard Cartesian set ${e_{x},e_{y},e_{z}}$ as $\mathbf{e}_{min}$, is the primary basis for computing the normal vector $\mathbf{n}_{GP}=R\mathbf{e}_{min}$. Using normals $A$ from the nearest point cloud points and Gaussian normals $B$, we calculate geometric loss with cosine similarity as follows:
$\mathcal{L}_{G}=1-\Bigg{|}\frac{A\cdot B}{||A||\text{ }||B||}\Bigg{|}$ (5)
@@
with $\mathcal{L}_{Image}$ an L1 loss and $\mathcal{L}_{D-SSIM}$, a structural similarity term, between rendered and ground truth images. This loss optimizes trade-offs between visual fidelity, depth error, and geometric alignment with empirically set weights.
SSIM ↑
MS-SSIM* —LPIPS ↓
Depth ↓
MSE
FPS↑
GPU ↓
min
ColonNerf[shi2023colonnerf] Phantom 25.54 0.86* 0.40 - - -
NeRF[mildenhall2020nerf] Simulation 35.29 0.92 0.14 0.007 $<$2 167
Phantom 32.10 0.81 0.39 4.263 $<$2 88
In-Vivo 18.93 0.67 0.43 0.109 $<$2 89
REIM- NeRF [psychogyios2023realistic] Simulation 32.22 0.82 0.33 $\bm{0.001}$ $<$2 170
Phantom 31.66 0.78 0.22 $\bm{0.013}$ $<$2 120
In-Vivo 18.94 0.65 0.45 $\bm{0.006}$ $<$2 97
Gaussian Pancakes (ours) Simulation $\bm{40.34}$ $\bm{0.97}$ $\bm{0.05}$ 0.007 $>$ $\bm{100}$ $\bm{0.83}$
Phantom $\bm{32.31}$ $\bm{0.90/1.00*}$ $\bm{0.20}$ 0.498 $>$ $\bm{100}$ $\bm{1.70}$
In-Vivo $\bm{26.25}$ $\bm{0.83}$ $\bm{0.21}$ 0.156 $>$ $\bm{100}$ $\bm{1.25}$
Table 2: Ablation study on In-Vivo data.
Method PSNR ↑
Depth ↑
SSIM
Depth ↓
MSE
GPU ↓
min
GS 26.116 0.444 0.158 1.211
GS+Depth 26.062 0.446 $\bm{0.145}$ $\bm{1.206}$
GS+Pancaking 26.211 $\bm{0.460}$ 0.154 1.329
GS+Pancaking+Depth(ours) $\bm{26.248}$ 0.458 0.156 1.249
@@ 3 Experiments
3.0.1 Dataset and Evaluation Metrics:
Our evaluation encompasses three distinct datasets: Simulation [zhang2020template], In-Vivo [ma2021rnnslam], and Phantom [bobrow2023colonoscopy]. The Simulation dataset, featuring sequences of the cecum, rectum, and sigmoid regions at a resolution of 320 $\times$ 240, is derived from software simulations in Unity using the graphics engine[zhang2020template] for image rendering. Depth and pose references, however, are obtained from RNNSLAM. The In-Vivo dataset comprises real colonoscopy sequences each at 270 $\times$ 216 resolution from the RNNSLAM study[ma2021rnnslam]. We selected three distinct sequences ("cecum_t4_b", "desc_t4_a", and "transt_t1_a") covering the cecum, descending colon and transcending sections from the public Phantom dataset (C3VD[bobrow2023colonoscopy]), with 1350 $\times$ 1080 resolution colonoscopy videos. C3VD’s simple camera trajectories caused large gaps in reconstructions due to insufficient scene coverage for RNNSLAM’s mesh surface fusion step, limiting us to choose three datasets with satisfactory RNNSLAM performance for evaluation.
@@ 3.0.1 Dataset and Evaluation Metrics:
Implementation: For a fair comparison with REIM-NeRF and NeRF we downsampled the resolution of the phantom dataset (C3VD) to 338 $\times$ 270, similar to what was tested in [psychogyios2023realistic]. We keep the original resolution of simulation and in-vivo datasets, since they are already close to [psychogyios2023realistic]. Our methodology initiates with RNNSLAM, rendering 10 FPS on an Intel i7 CPU and Nvidia GeForce GTX 1080 GPU. Gaussian Pancakes uses inferred depth maps for supervision, along with inferred camera poses and points for initial setup, across all training sequences. Subsequently, Gaussian Pancakes models trained following [kerbl20233d]’s hyperparameters, but are trained for 7,000 iterations instead of 30,000, densifying Gaussians until the 4,000th iteration. We set $\delta$ at 0.2 in Eq. 3, with hyperparameters in Eq. 6 $\lambda_{1},\lambda_{2},\text{and }\lambda_{3}$ finely tuned to 0.2, 0.6, and 0.2 (empirically set), respectively, balancing photometric, reconstruction, and geometric results. Geometric loss integration commences after the first 1000 iterations for non-uniform scaling. All models are trained for 1-2 minutes on an NVIDIA RTX A6000.
4 Results and Discussion
In Table 2.3, Gaussian Pancakes demonstrates performance improvements in photometric errors PSNR, SSIM, and LPIPS, and achieves over 100 FPS rendering speeds and training times less than 2% that of the other models for all datasets. Qualitative comparison is in the supplementary video from minute 3:46-4:43 and supplementary Fig. A.1, where it can be observed that NeRF and REIM-NeRF exhibit poorer performance on the In-Vivo dataset, even when compared to their own results on the other datasets, likely due to the reliance on ground truth camera poses and, for REIM-NeRF, accurate depth maps. The heightened noise in In-Vivo dataset’s camera poses and depth maps highlights Gaussian Pancakes’ superior ability to manage data noise. However, Gaussian Pancakes exhibited a higher Depth MSE across all datasets compared to REIM-NeRF, a result of using inferred depth maps for both training and evaluation; making Depth MSE a measure of residual training error rather than a comparison with actual ground truth. Our method effectively handles data noise, prioritizing high-quality images and smooth reconstructions even with suboptimal depth references. A comparison with ColonNeRF for all datasets was not possible due to unavailable code.
Ablation Analysis: Table 2 shows the incremental benefits of depth and geometric regularizations added to the foundational 3D GS method, initialized with RNNSLAM outputs, to reduce artifacts in unseen viewpoints. However, standard metrics fall short in capturing these improvements, where test views are too close to training views to reveal the artifacts. Nevertheless, the quantitative data reveals nuanced improvements: introducing depth regularization marginally reduces Depth MSE, and normal regularization (GS+Pancaking) improves Depth SSIM, suggesting better surface consistency. The combination of both (GS+Pancaking+Depth) achieves the best overall balance, as evidenced by modest gains in PSNR and Depth SSIM, alongside a negligible increase in processing time. Phantom and Simulation datasets show metric-based improvements that are more pronounced, further detailed in the Appendix Table A.1. For a comprehensive understanding, see the depth image comparison in supplementary Fig. A.3 and video from 4:43-5:40, which qualitatively shows artifact mitigation. Additionally, surfaces are visually smoother, seen in Fig. A.2.
5 Conclusion
Our evaluations across three varied datasets show that it achieves superior rendering quality, smoother reconstructions, fewer artifacts, and greatly reduces computational cost. Additionally, the method outputs explicit geometry, which sets it apart from the other leading methods with its potential for seamless integration into clinical practices. However, it’s important to note that the method’s success partly hinges on the quality of RNNSLAM’s surface fusion, facing occasional challenges in less consistent environments. Moving forward, we plan to integrate SuperPoint [barbed2022superpoint] for better point cloud creation and to refine depth map precision. This direction aims to not only bolster our method’s accuracy but also its applicability in a broader range of clinical scenarios.